OpenStudy (anonymous):

A delivery service charges a base price for an overnight delivery of a package plus an extra charge for each pound the package weighs. A customer is billed $22.85 for shipping a 3-pound package and $40 for shipping a 10-pound package. a.Write an equation that gives the total cost of shipping a package as a function of the weight of the package. b.Find the cost of shipping a 15-pound package.

OpenStudy (anonymous):

A) Let x be the cost per pound of the package and y be the fixed base cost of shipping. The two scenarios would have the following equations\[3x+y = 22.85\]\[10x+y=40\]if you subtract the two equations you can solve for x \[(10x-3x)+(y-y)=40-22.85\]\[7x=17.15\]\[x = 2.45\] So it costs $2.45 per pound to ship. Plug this into either equation and solve for y\[10(2.45)+y = 40\]\[y=15.5\]so it costs $15.50 just to ship the package. Thus a function for the cost of the package is \[f(x) = 2.45x + 15.50\] B) Plug 15 in for x in that equation to find the cost of shipping a 15 pound package

OpenStudy (anonymous):

is it 44.25

OpenStudy (anonymous):

52.25

OpenStudy (anonymous):

2.45*(15)+15.50